OpenStudy (didee):

Linear Algebra question: Suppose A is a 3 x 3 matrix such that det (A) = 4. Then det (4(A^-1)^T) is equal to?

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\]

OpenStudy (anonymous):

^T at the end means transpose correct? It's been quite a while since my linear algebra days.

OpenStudy (anonymous):

Well here is my reasoning ok so determinant of A =4 then \[\left| A^{-1} \right| = 1/\left| A \right|\] therefore if det(A)=4 then det(A^-1)=1/4. since we are multiplying by 4 then determinant is now 1 finally if i remember correctly... \[\left| A \right|=\left| A^T \right|\] there fore there answer should just be 1 i think. i would love for someone to double check my logic though if anyone else remembers this stuff.

OpenStudy (dumbcow):

looks good to me i have forgotten all the determinant properties

OpenStudy (anonymous):

Ok good thank you =). I had to reach WAY back for those ha ha

OpenStudy (anonymous):

i know the properties are right i was just unsure in the logic and how i followed them.

OpenStudy (dumbcow):

yeah i just looked up some determinant info you are correct :)

OpenStudy (anonymous):

NOOO i did it wrong. i got the right answer but i did my order of operations wrong i should of transposed first before i multiplied. so it should be det(A^-1)= 1/4 transpose = 1/4 mupltipy by 4 = 1

OpenStudy (didee):

thanks for the answer. So if i understand you first have to do the inverse A^-1 which gives you 1/A and then you have to get the transpose which gives?

OpenStudy (anonymous):

det(A)=det(A^T) so determinant of A is equal to the transposed determinant of A

OpenStudy (anonymous):

Just follow order of operations and use each property i gave where necessary

OpenStudy (didee):

Thanks, I have one more question. Can I post it here? I'm new to this.

OpenStudy (anonymous):

just post new thread its easier

OpenStudy (zarkon):

the answer is \[4^3\cdot\frac{1}{4}=16\]

OpenStudy (phi):

I see you beat me to it. But I made up this example in MATLAB, so I'll post it.

OpenStudy (didee):

thanks guys for your input